Please help me with this problem.
Find the length of LO if M is between L and O, LM=7x-9, MO=14 inches, and LO=10x-7. x=______ LO=________
I have x as being 4. But the "inches" are throwing me off.....I don't know where to go from that. I'm not even sure if I solved "x" correctly.
Thank You!

Answer by edjones(8007)

LM+MO=LO
7x-9+14=10x-7
7x+5=10x-7
7x-10x+5-5=10x-10x-7-5
-3x=-12
-3x/-3=-12/-3
x=4
10*4-7
=40-7
=33 in.
.
Ed